Given: LOGcX=3.1 LOGcY=1.8 LOGcZ=2.7
evaluate LOGc(XYZ)
evaluate LOGc(XY/Z)

Answer by lwsshak3(11628)

Given: LOGcX=3.1 LOGcY=1.8 LOGcZ=2.7
evaluate LOGc(XYZ)
evaluate LOGc(XY/Z)
..
logc(xyz)=logc(x)+logc(y)+logc(z) (using multiplication rule for logarithms)
=3.1+1.8+2.7=7.6
..
logc(xy/z)=logc(x)+logc(y)-logc(z) (using multiplication and division rules for logarithms)
=3.1+1.8-2.7=2.2